You are considering two salary offers. The first one will pay you $80,000 a year for two years. The second one will pay you $60,000 a year for two years and an $35,000 additional bonus, paid today.

The salaries would be paid in a lump sum at the end of each year. If the interest rate is 5%, compounded monthly, which is the better offer?  
Answer #1

Option1

Salary = $80000

Time = 2 years

Interest Rate = 5% compounded monthly

Effective Annual Interest Rate = (1+(0.05/12))12 -1 = 5.116%

Present Value of Salary = 80000/(1.05116)1 + 80000/(1.05116)2

Present Value of Salary = 76106.4 + 72402.3 = $148508.7

Option2

Salary = $60000

Time = 2 years

Additional Bonus paid today = $35000

Interest Rate = 5% compounded monthly

Effective Annual Interest Rate = (1+(0.05/12))12 -1 = 5.116%

Present Value of Salary = 60000/(1.05116)1 + 60000/(1.05116)2 + 35000

Present Value of Salary = 57079.8 + 54301.7 + 35000 = $146381.5

Answer Option 1 is better.